OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7.

OpenStudy (anonymous):

.y =1/28 x2 x = 1/28 y2 -28y = x2 y2 = 14x

OpenStudy (anonymous):

x = 1/28 y2

OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 2) and a directrix at y = -2. y2 = 2x y = 1/2 x2 y2 = 8x y =1/8 x2 @matricked

OpenStudy (anonymous):

y =1/8 x2

OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, 9). y =1/36 x2 y =1/9 x2 y2 = 9x y2 = 36x

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

y= =1/36 x2

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

i was offline and please don't specify me as others have more knowledge to sort the problems

OpenStudy (anonymous):

for Q5 its second option

OpenStudy (anonymous):

for Q6 let the eq of parabola be x^2=-k y now substiute (14,-74) or (-14,-74) to find k and then use the eq for ur question